### v2

# Mutual Information Estimation via Score-to-Fisher Bridge for Nonlinear Gaussian Noise Channels

Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoising score matching (DSM) learning for estimating the score function of channel output. Via de Bruijn’s identity, Fisher information estimated from the learned score function yields accurate estimates of MI through a Fisher integral representation for a variety of priors and channel nonlinearities. In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels.
I Introduction
We consider additive Gaussian noise channels with a deterministic nonlinear front-end,
$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (1)
and study how to estimate the mutual information (MI) $I(X;Y_{t})$ efficiently and accurately. A nonlinear front-end function can be represented by a deterministic function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$. Such models arise widely due to nonlinear front-end saturation/clipping, DAC/ADC nonlinearity, photonic modulators, optical fibers, or learned encoders implemented with neural networks. The value of the mutual information $I(X;Y_{t})$ is important for evaluating the ultimate performance of coded communication systems; it provides useful information for designing the system such as coding, preprocessing, or decoding/detection algorithms. In a sensing system, the nonlinear front-end can be interpreted as a details of sensing process. The value of the mutual information $I(X;Y_{t})$ indicates how much information is transferred from the sensing target to the output.
When $f$ is nonlinear, evaluating $I(X;Y_{t})$ is, in general, challenging in practice because accurate posterior means $\mathbb{E}[X|Y_{t}]$ and normalized likelihoods are inaccessible or computationally intractable. Let $W=f(X)$ be a transformed input. The de Bruijn identity immediately yields
$\frac{d}{dt}I(W;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (2)
so that the mutual information $I(X;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
Our key observation is that the Fisher information $J(Y_{t})$ can be estimated from the score function of $Y_{t}$ which can be estimated by denoising score matching (DSM) learning from samples of $Y_{t}$, and that the deterministic nonlinearity can be absorbed into a transformed input $W=f(X)$, so that
$I(X;Y_{t})=I(W;Y_{t}),\quad Y_{t}=W+Z_{t},$ (3)
holds for any deterministic function $f$. This equality is the basis of our score-based estimator of MI.
By integrating (2) over $t\in[T,\infty)$, for general $f$, we have the Fisher integral representation of MI:
$I(X;Y_{T})=\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt$ (4)
that depends only on the Fisher information of $Y_{t}$.
In this paper, we propose a score-based estimator of MI tailored to $Y_{t}=f(X)+Z_{t}$ [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi. The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4). Crucially, the approach never requires access to the posterior $p_{X|Y}(\cdot|\cdot)$. All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
The main contributions of this paper are summarized as follows:
  • Score-only MI estimation for nonlinear front-ends: We formalize a pipeline— DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be easily sampled.
  • Practical recipe: We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
  • Validation and use-cases: We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
In linear Gaussian channels, multivariate de Bruijn identity provides differential and integral representations of information quantities. For nonlinear $f$, however, posterior-free and normalization-free MI estimators remain limited. Building on DSM learning, our approach attaches directly to the marginal density $p_{Y_{t}}$, circumventing posterior computations and partition functions. Moreover, comparing with the conventional kernel density estimation (KDE) based estimators, our approach is robust to the high-dimensionality of the input space. The estimate of MI can be used for many applications such as design of coding systems, optimization of preprocessing, sensor design, and system optimization, and design of detection and decoding algorithms.
II Preliminaries
II-A Notation
Unless otherwise stated, all logarithms are natural (nats). Bold letters denote vectors in $\mathbb{R}^{n}$. For a random vector $U$ with density $p_{U}$, we write $h(U)$ for its differential entropy, $I(\cdot;\cdot)$ for mutual information, and $\operatorname{Cov}(U)$ for its covariance matrix. The (scalar) Fisher information is defined via the score function $s_{U}(\bm{u})\equiv\nabla_{\bm{u}}\log p_{U}(\bm{u})$ as
$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]$ (5)
see, e.g., [1, 2].
II-B Related works
For channels with memory, simulation-based methods have been developed to evaluate information rates without closed-form expressions. Arnold et al. [3] proposed a simulation-based framework that computes upper and lower bounds on the information rate by combining auxiliary finite-state channel models with forward–backward recursions and importance sampling; the method is broadly applicable to intersymbol interference (ISI) and finite-state channels where exact evaluation is infeasible. Building on this line, Sadeghi et al. [4] optimized these bounds over the choice of auxiliary models, yielding systematically improved estimates and a practical tool for design and benchmarking of channels with memory.
Another line of work estimates the MI directly from its definition via simulation-based plug-in estimators: one approximates $p_{Y}$ (or $h(Y)$) using KDE [5] or employs $k$-nearest neighbors (kkNN) estimators [6]. In the additive Gaussian setting, recent theory shows that plug-in estimation of $h(X+Z)$ is in fact optimal [7].
II-C AWGN channel and score function
Let $X\in\mathbb{R}^{n}$ be a channel input random vector with known prior probability density function (PDF) $p_{X}$, and consider the AWGN channel family:
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (6)
where $t$ represents the noise variance. The matrix $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by
$s_{Y_{t}}(\bm{y})\equiv\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y}).$ (7)
Throughout, we assume that sampling from $p_{X}$ is computationally feasible by forward sampling or Langevin sampling using the score $\nabla_{\bm{x}}\log p_{X}(\bm{x})$ (e.g., Rossky et al. [8], Welling and Teh [9]).
The Fisher information of $Y_{t}$ is given by
$J(Y_{t})\equiv\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\bigl[\|s_{Y_{t}}(\bm{y})\|^{2}\bigr],$ (8)
where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
$\mathbb{E}[X|Y_{t}=\bm{y}]=\bm{y}+ts_{Y_{t}}(\bm{y}).$ (9)
II-D I–MMSE relation
For the SNR-parameterized AWGN model
$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}),\quad N\perp X,$ (10)
the I–MMSE relation [13, 14] states that the identity
$\frac{d}{d\gamma}I(X;Y_{\gamma})=\frac{1}{2}\mathrm{mmse}(\gamma),$ (11)
holds for any snr $\gamma>0$ where $\mathrm{mmse}(\gamma)$ is defined by
$\mathrm{mmse}(\gamma)\equiv\mathbb{E}\bigl[\|X-\mathbb{E}[X|Y_{\gamma}]\|^{2}\bigr].$ (12)
With the change of variables $\gamma=1/t$ (and noting $I(X;Y_{\gamma})=I(X;Y_{t})$), an equivalent form for our variance parameterization $Y_{t}=X+Z_{t}$ is given by
$\frac{d}{dt}\,I(X;Y_{t})=-\frac{1}{2t^{2}}\mathrm{mmse}(t),$ (13)
where $\mathrm{mmse}(t)\equiv\mathbb{E}\bigl[\|X-\mathbb{E}[X|Y_{t}]\|^{2}\bigr].$
II-E de Bruijn identity
Let $h(Y_{t})$ denote the differential entropy of $Y_{t}$. Consider the AWGN channel setup defined by (6). Then the (multivariate) de Bruijn identity gives
$\frac{d}{dt}h(Y_{t})=\frac{1}{2}J(Y_{t}).$ (14)
The de Bruijn identity was discussed in [15, 16], which is proved by using the heat partial differential equation.
II-F Denoising score matching (DSM) method
DSM [17, 18, 19, 20] provides a practical way to learn the score function of a Gaussian-smoothed distribution without computing any normalization constants or partition functions. Let $p$ be a target density on $\mathbb{R}^{n}$ and define a Gaussian-smoothed sample:
$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}).$ (15)
Writing $p_{\sigma}\equiv p*\phi_{\sigma^{2}}$ for the convolution with a zero-mean Gaussian PDF $\phi_{\sigma^{2}}$ with variance $\sigma^{2}$, the DSM loss function for a parametric score model $s_{\bm{\theta}}:\mathbb{R}^{n}\!\to\!\mathbb{R}^{n}$ ($\bm{\theta}$ represents the vector of adjustable parameters) is given by
$\mathbb{E}_{\bm{w},\bm{\epsilon}}\Big[\big\|s_{\theta}(\bm{v})+\epsilon/\sigma\,\big\|^{2}\Big]$ (16)
$=\mathbb{E}_{\bm{v}\sim p_{\sigma}}\Big[\big\|\,s_{\theta}(\bm{v})-\nabla_{\bm{v}}\log p_{\sigma}(\bm{v})\,\big\|^{2}\Big]+\text{const},$ (17)
where this equivalence was shown by Vincent [18]. This means that minimizing the DSM loss function is equivalent to minimizing the expectation of the squared distance between the learned score function and the true score function, which is often called the Fisher divergence. This means that the unique minimizer of the DSM loss function satisfies $s_{\bm{\theta}}(\bm{y})\simeq\nabla_{\bm{y}}\log p_{\sigma}(\bm{y})$. Equivalently, since $\bm{v}-\bm{w}=\sigma\bm{\epsilon}$, one may implement the loss function as $(\bm{v}-\bm{w})/\sigma^{2}$:
$\mathbb{E}_{\bm{w},\bm{\epsilon}}\big[\|s_{\theta}(\bm{v})+(\bm{v}-\bm{w})/\sigma^{2}\,\|^{2}\big],$ (18)
which is an equivalent form of the DSM loss function.
In practice, we use a neural network to parameterize the score model $s_{\bm{\theta}}(\cdot)$. Such a score model is also called a score network. A mini-batch training with stochastic gradient descent or other variants such as Adam is typically used to minimize the DSM loss function.
III Estimating Mutual Information in AWGN Channels
In this section, we present a method to estimate the MI for AWGN channels described by
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (19)
where $X\sim p_{X}$. The results obtained in this section will be extended to general nonlinear channels in Section V.
III-A MI–Fisher relation
For any Gaussian channel, the following proposition holds.
Proposition 1.
For all $t>0$, the identity
$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (20)
holds [13, 14, 21].
Proof.
A sketch of proof is the following. Using the definition of MI,
$I(X;Y_{t})=h(Y_{t})-h(Z_{t})$ (21)
with $h(Z_{t})=(n/2)\log(2\pi e\,t)$, we have
$\frac{d}{dt}I(X;Y_{t})=\frac{1}{2}J(Y_{t})-\frac{n}{2t}$ (22)
by de Bruijn identity. ∎
It can be remarked that, from the proposition, we immediately obtain the MMSE–Fisher relation
$\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ (23)
by the I–MMSE relation.
Note that Reeves et al. [21] investigated related identities in the context of linear Gaussian channels, highlighting its implications for non-Gaussian input distributions. Their work extends the I–MMSE framework beyond the classical Gaussian setting.
III-B Fisher Integral representations of mutual information
From the identity (20) and the fact that $\lim_{t\to\infty}I(X;Y_{t})=0$ (since $Y_{t}$ and $X$ become independent in the limit), integrating over $t\in[T,\infty)$ yields
$I(X;Y_{T})=$ $\frac{1}{2}\int_{T}^{\infty}\!\Big(\frac{n}{t}-J(Y_{t})\Big)\,dt.$ (24)
The integrand decays as $O(1/t^{2})$, since
$0\ \leq\ \frac{n}{t}-J(Y_{t})\;=\;\frac{\mathrm{mmse}(t)}{t^{2}}\ \leq\ \frac{\operatorname{tr}\operatorname{Cov}(X)}{t^{2}}.$ (25)
The integral representation directly follows by combining the de Bruijn identity [15] with the I–MMSE relation [13, 14] as well.
III-C Estimation of Fisher Information
Using DSM learning, we can estimate the score function of $Y_{t}$. With $p\leftarrow p_{X}$ and $\sigma^{2}\leftarrow t$, the learned score model can approximate the score of $Y_{t}=X+Z_{t}$:
$s_{\bm{\theta}}(\bm{y})\approx\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=s_{Y_{t}}(\bm{y}),$ (26)
where $s_{\bm{\theta}}(\cdot)$ can be implemented with a neural network with parameter vector $\bm{\theta}$. This enables estimation of $J(Y_{t})=\mathbb{E}_{\bm{y}\sim p_{Y_{t}}(\cdot)}\|s_{Y_{t}}(\bm{y})\|^{2}$ from samples of $\bm{x}\sim p_{X}$ and Gaussian noises.
In particular, for a fixed noise level $t$, the DSM loss function can be written as
$\mathcal{L}_{t}(\bm{\theta})\;\equiv\;\mathbb{E}_{\bm{x},\bm{\epsilon}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})+\bm{\epsilon}/\sqrt{t}\,\right\|^{2}\right]$ (27)
$=\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]+\text{const},$ (28)
where the constant term does not depend on $\bm{\theta}$. Hence, with a sufficiently rich function class for $s_{\bm{\theta}}$ (e.g., a universal neural network) and sufficiently many samples, any sequence of empirical minimizers achieves $\mathbb{E}_{\bm{y}}\!\left[\left\|\,s_{\bm{\theta}}(\bm{y})-s_{Y_{t}}(\bm{y})\,\right\|^{2}\right]\rightarrow 0,$ i.e., the learned score converges to the true score in the Fisher divergence sense.
Moreover, with i.i.d. draws $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$ from $p_{Y_{t}}(i=1,2,...,N)$, the Monte Carlo estimator:
$\widehat{J}_{t}\triangleq\frac{1}{N}\sum_{i=1}^{N}\|s_{Y_{t}}(\bm{y}_{i})\|^{2}$ (29)
satisfies $\widehat{J}_{t}\xrightarrow{\text{a.s.}}J(Y_{t})$ as $N\to\infty$ (law of large numbers), and $\sqrt{N}\,(\widehat{J}_{t}-J(Y_{t}))\Rightarrow\mathcal{N}\big(0,\operatorname{Var}(\|s_{Y_{t}}(\bm{y})\|^{2})\big)$ (central limit theorem), so the estimation error decays at rate $O(N^{-1/2})$. In practice we plug in the learned score $s_{\bm{\theta}}$ in place of $s_{Y_{t}}$; under Fisher divergence consistency of $s_{\bm{\theta}}$, the plug-in estimator $(1/N)\sum_{i=1}^{N}\|s_{\theta}(\bm{y}_{i})\|^{2}$ is expected to be also consistent for $J(Y_{t})$.
III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$ where $P>0$ represents the signal power. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
Proposition 2 (Closed forms for Gaussian inputs).
Let $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
$\frac{1}{n}I(X;Y_{t})=\frac{1}{2}\log\!\Big(1+\frac{P}{t}\Big),$ (30)
$\frac{1}{n}h(X\mid Y_{t})=\frac{1}{2}\log\!\Big(2\pi e\,\frac{Pt}{P+t}\Big),$ (31)
$\mathrm{mmse}(t)=n\,\frac{Pt}{P+t},\quad J(Y_{t})=\frac{n}{P+t}.$ (32)
Sketch of proof: .
From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\mathbf{0},(P+t)\mathbf{I}_{n})$, hence we have
$s_{Y_{t}}(\bm{y})=\nabla_{\bm{y}}\log p_{Y_{t}}(\bm{y})=-(P+t)^{-1}\bm{y}$ (33)
and $J(Y_{t})=\mathbb{E}\|s_{Y_{t}}(Y_{t})\|^{2}=n/(P+t)$. The remaining identities follow by substituting $J(Y_{t})$ into $\mathrm{mmse}(t)=nt-t^{2}J(Y_{t})$ and by $h(X|Y_{t})=h(Y_{t})-I(X;Y_{t})$ with $h(Y_{t})=(n/2)\log\big((2\pi e)(P+t)\big).$
III-E Linear Gaussian channel case
Consider a linear Gaussian channel model:
$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (34)
where $A\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$. This setting preserves the Gaussian structure while introducing nontrivial correlation through the linear mixing due to $A$.
Proposition 3 (Closed forms for linear Gaussian channels).
Let $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$, $A\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n})$, independent. Then for all $t>0$:
$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}AA^{\top}\Bigr),$ (35)
$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}A^{\top}A\Bigr)^{-1}\right),$ (36)
$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((PAA^{\top}+tI_{n})^{-1}\Bigr).$ (37)
Sketch of proof:.
Since $AX$ is Gaussian with covariance $PAA^{\top}$, we have $Y_{t}\sim\mathcal{N}(\mathbf{0},PAA^{\top}+tI_{n})$. The mutual information (35) follows from the determinant formula for Gaussian differential entropy [2, 14]. The Fisher information (37) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(PAA^{\top}+tI_{n})^{-1}\bm{y}$. The MMSE formula (36) comes from the linear MMSE estimator of $X$ given $Y_{t}$. ∎
IV Estimation of Mutual Information
Using the score-based estimation of the Fisher information, we can evaluate $I(X;Y_{T})$ by numerical integration. In this section, we present details of the method to estimate the MI for the AWGN channel family described by (6).
IV-A Score-based Estimation Method for MI
The proposed estimation algorithm is summarized in Algorithm 1. The core of the algorithm is the DSM learning for $p_{Y_{t}}$ and the estimation of the Fisher information from the score model.
Algorithm 1 Score-to-Fisher bridge: Estimation of MI
1: Prepare a logarithmically spaced grid over $t$ (details of the log grid are described in Subsection IV-B).
2: Sample generation: Generate a random mini-batch $\mathcal{B}$ consisting of multiple pairs of $(\bm{x},\bm{\epsilon})$ for mini-batch training with forward sampling of $X$: $\bm{x}\sim p_{X},\ \bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n}).$
3: DSM learning for $p_{Y_{t}}$: train the score model parameters by minimizing the empirical DSM loss:
$\sum_{(\bm{x},\bm{\epsilon})\in\mathcal{B}}\big\|\,s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon})\;+\;\bm{\epsilon}/\sqrt{t}\,\big\|^{2}.$ (38)
Any stochastic optimizer, such as Adam, can be applied to minimize the empirical loss function.
4: Estimating the Fisher information: for each $t$, generate $\bm{y}_{1},\ldots,\bm{y}_{N}$ (e.g., by forward sampling $\bm{y}_{i}=\bm{x}_{i}+\sqrt{t}\,\bm{\epsilon}_{i}$), and compute the estimated Fisher information:
$\widehat{J}_{t}=\frac{1}{N}\sum_{i=1}^{N}\big\|s_{\bm{\theta}}(\bm{y}_{i})\big\|^{2}$ (39)
as an approximation of $J(Y_{t})$.
5: Estimating the MI via the Fisher integral representation:
$\widehat{I(X;Y_{T})}=\frac{1}{2}\int_{T}^{\infty}\Big(\frac{n}{t}-\widehat{J}_{t}\Big)\,dt,$ (40)
evaluated numerically.
In the above algorithm, we consider two alternatives for the DSM learning:
  • Scheme A (per-$t$ training). For each fixed $t$, minimize
    $\mathbb{E}\big\|s_{\theta}(\bm{x}+\sqrt{t}\bm{\epsilon})+\bm{\epsilon}/\sqrt{t}\big\|^{2}$
    to obtain the learned score model.
  • Scheme B (noise-conditional model). Learn $s_{\theta}(\bm{y},t)$ by minimizing
    $\mathbb{E}_{t\sim\mathcal{T}}\,\lambda(t)\ \mathbb{E}_{\bm{x},\bm{\epsilon}}\,\big\|s_{\bm{\theta}}(\bm{x}+\sqrt{t}\,\bm{\epsilon},t)+\bm{\epsilon}/\sqrt{t}\big\|^{2},$ (41)
    where $\mathcal{T}$ is a training distribution over $t$ (e.g., logarithmically spaced as described later) and $\lambda(t)$ is a weight function, e.g., $\lambda(t)=t$.
In the case of Scheme B, the score network has two inputs: the channel output $\bm{y}$ and the noise level $t$.
IV-B Log-domain trapezoid integration
The integrand in the Fisher integral representation (4) of the MI is smooth and may have a long tail. The choice of the integration grid is critical for achieving high numerical accuracy. In this subsection, we present a geometric (log-spaced) grid for the integration. For $t\in[t_{\min},t_{\max}]$ with $M$ points, take $\log t$ to be equally spaced:
$t_{k}=t_{\min}\left(\frac{t_{\max}}{t_{\min}}\right)^{\frac{k}{M-1}},\qquad k=0,1,\ldots,M-1.$ (42)
Equivalently in standard deviation, $\sigma_{k}\equiv\sqrt{t_{k}}$ with $\sigma_{k}=\sigma_{\min}(\sigma_{\max}/\sigma_{\min})^{k/(M-1)}$.
Rather than applying the trapezoid rule directly on the non-uniform $t$-grid, we perform the change of variables $u=\log t$ to obtain uniform spacing in the log domain. Using the integral representation (4), and substituting $u=\log t$ (so $dt=e^{u}du$ and $t=e^{u}$), we obtain
$I(X;Y_{T})=\frac{1}{2}\int_{\log T}^{\infty}\!\Big(n-e^{u}\,J(Y_{e^{u}})\Big)\,du.$ (43)
Since the $t$-grid (42) corresponds to uniform spacing in $u=\log t$ with step size $\Delta_{u}\equiv\log(t_{\max}/t_{\min})/(M-1)$, we apply the standard trapezoid rule with uniform spacing:
$\widehat{I}(X;Y_{T})\approx\frac{1}{2}\sum_{k=k_{T}}^{M-2}\frac{\ell(u_{k})+\ell(u_{k+1})}{2}\,\Delta_{u}\;+\;\text{tail},$ (44)
where $\ell(u)\equiv n-e^{u}\,\widehat{J}(Y_{e^{u}})$, and $k_{T}$ is the smallest index with $t_{k_{T}}\geq T$. This log-domain integration offers several benefits over direct $t$-domain trapezoid integration: (1) uniform grid spacing reduces discretization error, (2) the transformation naturally handles a wide dynamic range in $t$, and (3) convexity bias is significantly reduced. In our experiments, this approach improved MI estimation accuracy by over an order of magnitude.
The tail contribution above $t_{\max}$ can be estimated using the asymptotic behavior. For Gaussian inputs, $\mathrm{mmse}(t)\to\operatorname{tr}\mathrm{Cov}(X)$ as $t\to\infty$, giving
$\text{tail}\approx\frac{1}{2}\int_{t_{\max}}^{\infty}\frac{\mathrm{mmse}(t)}{t^{2}}\,dt\approx\frac{1}{2}\,\frac{\operatorname{tr}\mathrm{Cov}(X)}{t_{\max}}.$ (45)
The trace $\operatorname{tr}\mathrm{Cov}(X)$ can be estimated from samples of $p_{X}$ and used as a tail correction.
While the tail correction formula (45) is exact for Gaussian inputs, its leading-order term $\mathrm{tr}\,\mathrm{Cov}(X)/t_{\max}$ remains valid for arbitrary inputs, since $\mathrm{mmse}(t)\to\mathrm{tr}\,\mathrm{Cov}(X)$ as $t\to\infty$. Therefore, the same tail correction can be applied as a universal approximation, with the understanding that higher-order terms may depend on the non-Gaussianity of $X$.
V Nonlinear channel
Let us consider our primal target channel:
$Y_{t}=f(X)+Z_{t},$
where $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{n}$ is a deterministic function [^2] An extension to a general case $f:\mathbb{R}^{n}\rightarrow\mathbb{R}^{m}$ is straightforward but we assume $m=n$ for simplicity.. The mutual information equivalence presented in the following proposition is a key observation which enables the score-based estimation to be applied to general nonlinear channels.
V-A Mutual information equivalence
Let $W\equiv f(X)$ be a measurable deterministic mapping, and consider $Y_{t}=W+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$. Then the conditional law of $Y_{t}$ given $X$ depends on $X$ only through $W$. This means that $X\to W\to Y_{t}$ forms a Markov chain and equivalently $Y_{t}\perp X\mid W$.
We can prove mutual information equivalence in the following proposition.
Proposition 4.
Let $Y_{t}=W+Z_{t}(t>0)$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$ independent of $(X,W)$ and $W=f(X)$, where $f$ is a measurable deterministic function. Then we have
$I(X;Y_{t})=I(W;Y_{t}).$ (46)
Proof.
By the chain rule of mutual information, we have
$I(X;Y_{t})=I(W;Y_{t})+I(X;Y_{t}\mid W)-I(W;Y_{t}\mid X).$ (47)
Here $I(X;Y_{t}\mid W)=0$ because of the conditional independence $Y_{t}\perp X\mid W$, and $I(W;Y_{t}\mid X)=0$ holds because $W$ is a function of $X$. Hence $I(X;Y_{t})=I(W;Y_{t})$. An equivalent proof uses the data processing inequality twice: $X\to W\to Y_{t}$ gives $I(X;Y_{t})\leq I(W;Y_{t})$, while $W\to X\to Y_{t}$ (since $W=f(X)$) gives $I(W;Y_{t})\leq I(X;Y_{t})$, yielding equality. ∎
From Proposition 4, we have the MI equivalence $I(X;Y_{t})=I(W;Y_{t})$, which implies that the score-based estimation method presented in the previous section can be applied to evaluate $I(X;Y_{T})$. Namely, instead of $X$, from the samples of $W$, we can estimate the Fisher information of $Y_{t}$ and the estimate of the Fisher information can be used to evaluate $I(W;Y_{t})$. The equivalence relation $I(X;Y_{t})=I(W;Y_{t})$ enables to obtain the estimate of $I(X;Y_{t})$ from the estimate of $I(W;Y_{t})$. Note that no invertibility of $f$ is required in this estimation method.
V-B Score-based estimation method for nonlinear channels
In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 should be replaced by the following steps:
  • Sample $W$ by $W=f(X)$ where $X\sim p_{X}$.
  • Estimate the Fisher information of $Y_{t}$ from the samples of $W$ by the score-based estimation in Algorithm 1.
  • Use the estimated Fisher information to evaluate the Fisher integral form of MI.
  • Use the equality $I(X;Y_{t})=I(W;Y_{t})$ to obtain the estimate of $I(X;Y_{t})$.
The estimation procedure is still a single-loop procedure, which does not require the access to the posterior density $p_{X|Y_{t}}(\cdot|\cdot)$.
VI Numerical experiments
VI-A Gaussian input case for validation
To validate the proposed score-based estimation method, we first consider the Gaussian input case where exact closed-form expressions are available. Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(0,PI_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(0,I_{n})$.
VI-A1 Experimental Setup
We implemented the DSM learning for $p_{Y_{t}}$ with the following configuration: The score network is implemented as a fully connected neural network consisting of 3 hidden layers, each containing 128 units with SiLU (Swish) activation functions. The input and output dimensions match the problem dimension $n$, where we test with $n\in\{4,8,16\}$ to evaluate scalability. For the training configuration, we set the signal power to $P=1.0$ and construct a logarithmically spaced grid of $M=10$ noise variance values ranging from $t\in[P/200,200P]$, which covers four orders of magnitude. At each noise level $t$, we train a separate score network for 300 iterations using mini-batches of 4096 samples. The optimization is performed using the Adam optimizer with a learning rate of $10^{-3}$, and gradient clipping is applied with a maximum norm of 1.0 to ensure stable training.
VI-A2 Validating the Fisher Information Estimate
For Fisher information estimation, we generate 100,000 Monte Carlo samples at each noise variance $t$ to ensure statistical reliability. The ground truth values are computed using the analytical formula $J(Y_{t})=n/(P+t)$, which is available for Gaussian input distributions.
Figure 1 presents the estimated Fisher information $\hat{J}(Y_{t})$ compared with the analytical ground truth $J(Y_{t})=n/(P+t)$ for dimensions $n\in\{4,8,16\}$. The results demonstrate excellent agreement between the DSM-based estimates and theoretical values across multiple orders of magnitude in the noise variance $t$.
Figure 1: Fisher information $J(Y_{t})$ versus noise variance $t$ for Gaussian input with $P=1$ and dimensions $n\in\{4,8,16\}$. Solid lines with circular markers show ground truth values $J(Y_{t})=n/(P+t)$, while dashed lines with triangular markers show estimated values $\hat{J}(Y_{t})$.
The results confirm that:
  1. 1.
    The DSM-based score estimation accurately captures the Fisher information across a wide range of noise levels ($t\in[0.005,200]$).
  2. 2.
    The method scales well with dimension, maintaining low relative errors even for $n=16$.
  3. 3.
    The median relative error remains below 1.3% for all tested dimensions, validating the theoretical framework.
VI-A3 Mutual Information Estimation
We demonstrate the complete pipeline from the Score-to-Fisher bridge for MI estimation using the integral representation of MI. The experiment uses the same configuration as previous tests with $n=4$ and $P=1.0$, implementing log-domain trapezoid integration with tail correction. Figure 2 presents the estimated mutual information (per symbol) $\hat{I}(X;Y_{t})/n$ compared with the analytical ground truth $I(X;Y_{t})/n=(1/2)\log(1+P/t)$ for Gaussian input. The DSM-based estimates demonstrate excellent agreement with the theoretical values across the entire noise variance range.
Figure 2: Mutual information (per symbol) $I(X;Y_{t})/n$ versus noise variance $t$ for Gaussian input.
A key contribution of this experiment is the log-domain trapezoid integration method. Compared to standard trapezoid integration on the $t$-grid, the log-domain approach improved estimation accuracy by over an order of magnitude, reducing median relative errors from approximately 26% to 0.4%. This substantial improvement arises from: (1) uniform grid spacing in the log domain, (2) reduced convexity bias, and (3) better numerical handling of the wide dynamic range in $t$ values.
The tail correction, based on the asymptotic behavior $\text{mmse}(t)\sim\text{tr Cov}(X)$ for large $t$, proves essential for maintaining accuracy across the full integration range. Without tail correction, the 90th percentile error degrades significantly to over 40%.
These validation results on Gaussian inputs, where analytical solutions exist, provide strong evidence for the reliability of the proposed Score-to-Fisher bridge method and related theoretical framework presented in the previous sections.
VI-A4 Noise-conditional model
We investigate a noise-conditional score model $s_{\bm{\theta}}(\bm{y},t)$ (Scheme B) that takes both observation $\bm{y}$ and noise variance $t$ as inputs, eliminating the need for separate per-$t$ training. The model uses Gaussian Fourier projection to embed scalar $t$ values, which are injected into each hidden layer. Training samples $t$ from a log-uniform distribution over $Pt_{\min},Pt_{\max}]$ with loss weighting $\lambda(t)=t$ for 20,000 steps.
For Gaussian input with $n=4$ and $P=1.0$, this approach achieves 6.61% median relative error across 12 evaluation points spanning $t\in[0.005,50]$. Figure 3 demonstrates reasonable agreement with analytical values across four orders of magnitude in noise variance. We can observe the tendency that the noise-conditional DSM estimate slightly underestimates the MI at low noise variance. This result may be due to insufficient training iterations for a conditional model. However, this unified framework requires only one neural network instead of 12 separate models, significantly reducing computational cost.
Figure 3: Noise-conditional model for mutual information estimation results for Gaussian input ($n=4$, $P=1$). Scheme B was used.
VI-B Discrete input case
VI-B1 Mutual Information Estimation
To validate the approach for discrete input distributions, we consider binary phase shift keying (BPSK) modulation with $X\in\{-\sqrt{P},+\sqrt{P}\}$ (equal probability), $P=1.0$, and dimension $n=1$. Unlike continuous distributions, BPSK admits exact mutual information calculation via numerical integration of the Gaussian mixture distribution providing ground truth for validation.
We implement per-$t$ training (Scheme A) over $M=12$ logarithmically spaced points in $t\in[0.005,50]$, with 1000 training iterations per noise level and 200,000 Monte Carlo samples for Fisher information estimation.
Figure 4 compares DSM estimates with exact BPSK mutual information and the Gaussian input baseline $I_{\text{Gauss}}(X;Y_{t})/n=\frac{1}{2}\log(1+P/t)$. The DSM estimates achieve excellent agreement with exact values, with median relative errors of 0.6% for low noise ($t<0.1$) and 2.8% for mid-range ($0.1\leq t<5$).
Figure 4: Mutual information estimation for BPSK input ($n=1$, $P=1$). Dotted line: Gaussian input theoretical baseline; solid circles: BPSK exact MI via numerical integration; dashed squares: DSM estimates.
VI-C Linear Gaussian channel case
To validate the method beyond the standard AWGN case, we consider the linear Gaussian channel:
$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (48)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.
We train separate score networks (3 hidden layers, 128 units each) for $M=10$ logarithmically spaced noise levels in $t\in[P/200,50P]$. Each network is trained for 300 iterations with 8192-sample mini-batches, and Fisher information is estimated using 100,000 Monte Carlo samples per noise level.
Figure 5 compares the DSM-based estimates with the analytical solution (per symbol) $I(X;Y_{t})/n=(1/2)\log\det(I_{n}+(P/t)AA^{\top})$. The results demonstrate excellent agreement across 4 orders of magnitude in noise variance, confirming that the Score-to-Fisher bridge method generalizes successfully to linear Gaussian channels with nontrivial correlation structure.
Figure 5: Mutual information estimation for linear Gaussian channel $Y_{t}={\bm{A}}X+Z_{t}$ ($n=4$, $P=1$). Circles: analytical ground truth; triangles: DSM estimates.
VI-D Nonlinear channel case
To demonstrate the effectiveness of the proposed method for nonlinear channels where closed-form solutions are unavailable, we consider a composite channel with linear mixing followed by elementwise tanh nonlinearity:
$Y_{t}=\tanh({\bm{A}}X)+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (49)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $\tanh(\cdot)$ is applied elementwise. This model represents practical scenarios where signals undergo linear transformations before nonlinear processing.
VI-D1 Experimental Setup
We use the same DSM configuration as in the linear case: $n=4$, $P=1.0$, with $M=12$ logarithmically spaced noise levels in $t\in[P/200,50P]$. The orthogonal matrix $\bm{A}$} is generated via QR decomposition of a random Gaussian matrix. For each $t$, we train score networks for 400 iterations using 8192-sample mini-batches. Fisher information estimation uses 100,000 Monte Carlo samples per noise level.
VI-D2 KDE-LOO Baseline
Since analytical solutions are unavailable for nonlinear channels, we implement a kernel density estimation with leave-one-out (KDE-LOO) baseline [6] that directly estimates mutual information from its definition. For samples $\{(\bm{w}_{j},\bm{y}_{j})\}_{j=1}^{N}$ with $\bm{y}_{j}=\bm{w}_{j}+\sqrt{t}\bm{\epsilon}_{j}$, the KDE-LOO estimator is defined by
$\hat{I}_{\text{KDE}}(X;Y_{t})=\frac{1}{N}\sum_{i=1}^{N}\Bigg[$ $-\frac{\|\bm{y}_{i}-\bm{w}_{i}\|^{2}}{2t}$
$-\log\left(\frac{1}{N-1}\sum_{j\neq i}e^{-\frac{\|\bm{y}_{i}-\bm{w}_{j}\|^{2}}{2t}}\right)\Bigg].$ (50)
This estimator is derived from the plug-in principle applied to $I(X;Y_{t})=\mathbb{E}[\log p_{Y_{t}|X}(\bm{y}|\bm{x})]-\mathbb{E}[\log p_{Y_{t}}(\bm{y})]$, where the marginal density $p_{Y_{t}}(\bm{y})$ is approximated using LOO kernel density estimation to avoid overfitting. We use $N=20,000$ samples with full-sum computation (no $k$-nearest neighbor approximation) to ensure high accuracy.
VI-D3 Results and Discussion
Figure 6 presents the comparison between DSM-based estimates and the KDE-LOO baseline for the composite nonlinear channel. The results demonstrate remarkable agreement between the two methods across the entire noise variance range, with both curves nearly overlapping.
Figure 6: Mutual information estimation for composite nonlinear channel $Y_{t}=\tanh({\bm{A}}X)+Z_{t}$ ($n=4$, $P=1$). Circles: KDE-LOO baseline; triangles: DSM estimates.
Key observations include:
  1. 1.
    Excellent agreement: The DSM estimates closely match the KDE-LOO baseline across four orders of magnitude in noise variance, with relative errors typically below 5%.
  2. 2.
    Computational efficiency: While KDE-LOO requires $O(N^{2})$ distance computations per noise level, DSM achieves comparable accuracy with $O(N)$ score evaluations after one-time training where $N$ is the number of samples.
  3. 3.
    Nonlinear validation: This result provides strong evidence for the Score-to-Fisher bridge approach in scenarios where analytical benchmarks are unavailable.
The success of this experiment confirms that the proposed method generalizes effectively to handle composite channels combining linear mixing and nonlinear transformations, making it a practical tool for mutual information estimation in realistic communication systems.
VII Conclusion
We have presented a novel approach for mutual information estimation in nonlinear Gaussian noise channels via the Score-to-Fisher bridge approach. The method leverages DSM learning of score model for $p_{Y_{t}}$ to estimate Fisher information, which is then converted to mutual information via the Fisher integral representation.
Experimental validation across linear Gaussian channels, composite nonlinear channels, demonstrates excellent agreement with analytical solutions and baseline methods. The approach achieves computational efficiency with $O(N)$ evaluations per noise level while maintaining high accuracy across wide dynamic ranges of the noise variance. The curse of dimensionality can be avoided due to the efficiency of the score model evaluation.
The Score-to-Fisher bridge methodology provides a practical and theoretically sound framework for MI estimation in realistic nonlinear Gaussian noise channel models where closed-form solutions are unavailable. Accurate estimation of MI plays a crucial role in assessing the fundamental limits of coded communication systems, and also serves as a valuable guideline for system design choices such as coding schemes, preprocessing strategies, and detection or decoding algorithms.
Acknowledgments
This work was supported by JST, CRONOS, Japan Grant Number JPMJCS25N5.